OpenStudy (anonymous):

Approximate the real number solutions to the polynomial function f(x) = x3 – 4x2 – 4x + 16

hero (hero):

f(x) = x^3 - 4x^2 -4x + 16 f(x) = x^2(x-4) -4(x-4) f(x) = (x-4)(x^2-4) 0 = (x-4)(x^2-4) x - 4 = 0 x^2 - 4 = 0 x = 4 x^2 = 4 x = ±2 Therefore, 3 solutions